Solve by elimination...
3x-y=21
2x+y=4

Answer by Thanenjayan(7)

3x-y=21 ----- 1
2x+y=4 ------ 2

(1)+(2)=> 5x = 25
x = 5

substitude x=5 in (2)
2*(5)+y = 4
y = 4-10
y = -6